OpenStudy (anonymous):

The area of a rectangular wall of a barn is 56 square feet. Its length is 6 feet longer than twice its width. Find the length and width of the wall of the barn.

OpenStudy (anonymous):

Let us assume width is x feet therefore length = 2x +6 Area = x*(2x+6) 56 = 2x^2 + 6x x^2 + 3x - 28 = 0 x^2 + 7x - 4x - 28 = 0 x(x+7)-4(x+7) = 0 (x-4)(x+7) = 0 Either x = 4 or, x = -7(but x cannot be negative) So width = 4 feet and length = 2x + 6 = 2*4 + 6 = 14 feet